### v1

# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

Abstract
Food production is a vital global concern and the potential for an agritech revolution through artificial intelligence (AI) remains largely unexplored. This paper presents a comprehensive review focused on the application of machine learning (ML) in agriculture, aiming to explore its transformative potential in farming practices and efficiency enhancement. To understand the extent of research activity in this field, statistical data have been gathered, revealing a substantial growth trend in recent years. This indicates that it stands out as one of the most dynamic and vibrant research domains. By introducing the concept of ML and delving into the realm of smart agriculture, including Precision Agriculture, Smart Farming, Digital Agriculture, and Agriculture 4.0, we investigate how AI can optimize crop output and minimize environmental impact. We highlight the capacity of ML to analyze and classify agricultural data, providing examples of improved productivity and profitability on farms. Furthermore, we discuss prominent ML models and their unique features that have shown promising results in agricultural applications. Through a systematic review of the literature, this paper addresses the existing literature gap on AI in agriculture and offers valuable information to newcomers and researchers. By shedding light on unexplored areas within this emerging field, our objective is to facilitate a deeper understanding of the significant contributions and potential of AI in agriculture, ultimately benefiting the research community.
keywords:
Agriculture , Artificial Intelligence (AI) , Machine Learning (ML) , Precision Agriculture , Smart Farming , Digital Agriculture
1 Introduction
“The future of agriculture is not about better farming, but about smarter farming.” - Unknown
The dawn of human civilization can be traced back to the advent of agriculture, which has had an indelible impact on humanity. Agriculture encompasses a wide range of practices, including crop production [1] and livestock production, aquaculture, and forestry, which provide food and non-food products such as fibers, fuels, and raw materials such as rubber. It is an essential sector for feeding the growing world population, which is predicted by the Food and Agriculture Organization (FAO) to reach approximately nine billion by 2050 [2]. Agriculture also plays a critical role in the global economy, providing a foundation for sustainable growth and long-term economic development. Many countries have recognized the importance of this sector, as demonstrated by its contribution (Fig. 1) to their gross domestic product (GDP). In 2022 alone, the United States’ agricultural GDP reached \$201 billion in the fourth quarter of 2022, while China, Japan, and India also made notable contributions, reflecting the economic weight of the sector. The agricultural exports of the top five countries are depicted in Fig. 2. The increasing demand for food requires [3] continued focus on agricultural innovation and investment. To assess the global distribution of the agricultural land area, data was collected in the top $10$ countries ranked by agricultural land area in 2020 (Fig. 3). These countries not only demonstrate a significant commitment to agricultural production, but also play a crucial role in the global food supply chain.
Figure 1: GDP of top five counties [4].
Throughout history, three significant agricultural revolutions have shaped agricultural practices [5]. To enhance comprehension, a visual representation is provided in Fig. 4. The Neolithic Revolution, also known as the First Agricultural Revolution, emerged around 10,000 BCE in the Middle East’s Fertile Crescent. It witnessed the domestication of plants and animals and the rise of agriculture as a way of life. The Second Agricultural Revolution, which occurred in the $18^{th}$ and $19^{th}$ centuries in Europe and North America, introduced technological advances such as plow, seed drilling, and crop rotation, significantly increasing productivity. The Third Agricultural Revolution, known as the Green Revolution, was a global phenomenon of the mid-$20^{th}$ century. It involved the development of high-yielding crop varieties, the increase in the use of fertilizers and pesticides, the mechanization and irrigation, all of which aimed to improve agricultural productivity. Now, the fourth agricultural revolution, Agriculture $4.0$, is underway, with digital advancements and AI [6] playing a crucial role, akin to their impact on the fourth industrial revolution according to experts [7, 8].
Figure 2: Export in agriculture of top five counties [9] [10] [11].
ML is an interdisciplinary field that merges computer science and statistics to perform tasks that are typically performed by humans. It relies on data analysis and classification to train machines without explicit programming. In recent years, the integration of ML with big data technologies and high-performance computing has gained popularity in agriculture, facilitating the analysis of data-intensive processes. ML has become an essential tool in agriculture and has found applications in various scientific fields such as bio-informatics [12], biochemistry [13], medicine [14], meteorology [15, 16], economic sciences [17, 18], robotics [19], aquaculture [20], food security [21], optimization [22, 23], industry [24, 25], and climatology [26].
Figure 3: Agricultural land world wise, sq. km., 2020 - Country rankings [27].
Agriculture 4.0 represents a new era of farming, merging traditional methods with technologies such as IoT, AI, and automation. Through the use of sensors, drones, and AI-driven machinery, this approach increases efficiency and sustainability, enhancing yields, and reducing waste. It is a direct response to growing food needs and environmental challenges, reflecting the agricultural sector’s readiness to embrace innovation and adapt to a rapidly changing world. Precision agriculture [28], digital agriculture [29], and smart agriculture [30] are terms used to describe the use of technology and data in agricultural practices. These evolving agricultural practices hold great promise for improving efficiency, sustainability, and productivity in agriculture. Although they share some similarities, there are subtle differences in their focus and approach as follows:
  • Precision agriculture: This primarily aims to optimize the use of inputs such as water, fertilizers, and pesticides to improve crop yields.
  • Digital agriculture: This focuses on using digital technologies such as sensors, satellite imagery, and data analytics to collect and analyze data for better decision-making.
  • Smart agriculture: This involves the use of advanced technologies such as AI, robotics, and the IoT to automate and optimize agricultural processes.
Figure 4: Revolution in agriculture and key characteristics.
However, the agriculture sector faces challenges that include climate change, limited resources, and the need for sustainable practices. Unpredictable weather patterns, pests, and diseases affect crop yields. Limited land, water and energy resources strain production. ML offers innovative solutions to improve productivity and address these challenges. Although some papers have explored various ML models, a comprehensive review is still needed in this emerging field. This paper aims to provide that review, focusing on the unique features and applications of recent ML models in agriculture.
The organization of the paper is as follows: Section 2 begins by outlining the scope of the studies covered in this article. Section 3 serves as an overview of ML. Section 4 shifts the focus to agriculture, explaining its various steps. Section 5 reviews recent applications of ML in different stages of farming. A quantitative analysis is presented in Section 6. Few case studies are discussed in Section 7. The transformative influence of the top $10$ AI and robotics companies in agriculture is explored in Section 8. Section 9 delves into the challenges associated with the use of AI and ML in agriculture. Looking ahead, Section 10 presents future expectations within this field. The paper comes to its conclusion in Section 11, combining the key findings and implications of the study.
2 Scope of Studies
The field of agricultural AI has seen a significant surge in research, resulting in numerous reviews and state-of-the-art research articles published over time. In Table 1, we provide a summary of some of these review articles, highlighting their focused areas and key comments. In this review article, our aim is to address the gaps observed in the existing literature to offer the best understanding of the current research landscape in agricultural AI. Our methodology consisted of a three-stage process: identifying relevant studies, screening them, and evaluating their eligibility. This approach enabled us to outline the distinct features, advantages, and shortcomings of each application, thus illuminating both the potential benefits and challenges inherent in the deployment of ML in agricultural settings. Unlike previous surveys, this review offers a comprehensive look at practical agricultural applications of top companies $10$ in AI and robotics or start-ups in agriculture and how academic research can drive innovation.
This paper aims to explore the following key areas:
  1. RO1:
    Assessing the effects and results of applying AI and ML technologies in agriculture.
  2. RO2:
    Investigating the latest developments and research in AI technologies within the agricultural industry.
  3. RO3:
    Gaps between academic research and practical industry implementation of AI in agriculture.
Table 1: Summary of Review Papers on AI in Agriculture.
Analysis Type of Review
Ref. Pesticide Mngmnt Weed Detect Diseases Detect Crop Recmd. Sys. Auto. Irrig. Control Yield Mngmnt Soil/Seed Systematic Bibliogr. Comprehen. Mixed Rev.
[31]
[32]
[33]
[34]
[35]
[36]
[37]
[31]
This paper
However, our review paper does not focus on the technical intricacies of individual ML algorithms, nor does it extensively explore the broader socio-economic impacts of agricultural AI. Instead, our objective is to present a holistic and valuable analysis that offers readers a comprehensive understanding of the current state and future potential of ML in the transformation of agriculture.
3 Machine Learning: An Overview
ML involves a learning process in which algorithms acquire knowledge from training data to perform tasks. It is a field of AI focused on developing algorithms that learn from data to make predictions or decisions. By training models with data, ML enables accurate predictions or decisions with new data. Pioneered by Arthur Samuel [38] and Frank Rosenblatt in the mid-twentieth century, ML has advanced with the availability of powerful computers and extensive datasets. With the exponential growth of data in recent years, ML has become invaluable for organizations to make informed decisions, improve products and services, and optimize operations. ML finds applications in diverse industries such as finance, healthcare, transportation, and entertainment [39]. The three main types of ML are shown in Fig. 5.
Figure 5: Major classification of ML.
3.1 Supervised learning
Supervised learning is a type of ML in which a model is trained on a labeled data set, which means that the data set includes both input data and the corresponding output labels [40]. The objective of supervised learning is for the model to learn to map inputs to outputs accurately so that it can make accurate predictions or classifications when presented with new unseen data. The training process involves using various algorithms to analyze the labeled data set and develop a model that can accurately generalize and make predictions on new data. The details and key features of popular supervised learning models, including Linear Regression, Random Forests, Support Vector Machines, and Artificial Neural Networks, are presented in Table 2.
Table 2: Comparison of supervised learning Models.
SL Model Key Features Pros Cons
1 Linear Regression Simple to implement, easy to interpret Fast, works well with linearly separable data Limited to linear relationships
2 Logistic Regression Can handle binary and multi-class classification, interpretable coefficients Fast, works well with linearly separable data Limited to linear decision boundaries
3 Decision Trees Non-parametric, can handle both categorical and continuous data, easy to interpret Can capture complex non-linear relationships, robust to outliers Can overfit to the training data
4 Random Forests Ensemble of decision trees, reduces over-fitting, handles high-dimensional data Can handle complex non-linear relationships, robust to outliers Computationally expensive
5 Support Vector Machines (SVMs) Can handle high-dimensional data, works well with non-linearly separable data Can capture complex non-linear relationships Computationally expensive, sensitive to parameter settings
6 Naive Bayes Classifiers Simple and fast, can handle high-dimensional data Requires strong independence assumptions, may not work well with correlated features Limited to linear decision boundaries
7 Artificial Neural Networks (ANNs) Can learn complex non-linear relationships, can handle large amounts of data Can handle complex non-linear relationships Computationally expensive, requires careful tuning
8 K-Nearest Neighbor (KNN) Classifiers Simple and fast, non-parametric, can handle multi-class classification Can handle complex non-linear relationships, robust to outliers Computationally expensive at test time
9 Gradient Boosting Machines (GBMs) Ensemble of decision trees, handles heterogeneous data, reduces bias Can handle complex non-linear relationships Computationally expensive, sensitive to parameter settings
10 Convolutional Neural Networks (CNNs) Designed for image data, can learn hierarchical representations State-of-the-art performance on image classification tasks Computationally expensive, requires large amounts of data
3.2 Unsupervised learning
Unsupervised learning is an ML technique in which the model is trained on a data set without any predefined labels or target variable [40]. The goal is to identify patterns and relationships within the data, such as clusters or groups, and discover meaningful insights from the data. Unsupervised learning algorithms do not have a clear objective or specific output to predict, but rather aim to discover the underlying structure and relationships within the data. Examples of unsupervised learning algorithms include clustering, dimensionality reduction, and anomaly detection. The key features, pros, and cons of the most popular models in unsupervised learning are summarized in Table 3.
Table 3: Comparison of unsupervised learning models.
SL Model Key Features Pros Cons
1 Clustering Groups data points based on similarity, useful for exploratory data analysis Can identify unknown patterns, useful for outlier detection Can be sensitive to initialization, difficult to interpret results
2 Dimensionality Reduction Reduces the number of features in data, useful for visualization and data compression Can improve model performance by removing noise and irrelevant features May lose important information, difficult to interpret results
3 Association Rule Learning Identifies relationships between items in large datasets, useful for market basket analysis Can uncover hidden patterns and associations Limited to categorical data, may produce spurious associations
4 Anomaly Detection Identifies unusual or rare data points, useful for fraud detection Can identify unknown patterns and anomalies Difficult to determine appropriate threshold for anomaly detection
5 Generative Models Learns the underlying probability distribution of data, useful for data generation and synthesis Can generate new samples of data, useful for data augmentation Computationally expensive, difficult to train and interpret
3.3 Reinforcement learning
Reinforcement learning can be described as a category of ML in which an agent acquires the ability to act within an environment by taking specific actions and receiving rewards or punishments in response [40]. The goal of the agent is to learn to make the best decisions possible in order to maximize its rewards over time. The agent receives feedback from the environment in the form of rewards, which are positive or negative signals indicating whether the agent’s actions were good or bad. As the agent continues to interact with its environment, it gradually gains knowledge of which actions result in the greatest rewards and which should be avoided. This iterative learning process is commonly employed in a variety of applications, including gaming, robotics, and decision-making in intricate settings. Table 4 summarizes the key features, pros, and cons of the most popular models in reinforcement learning.
Table 4: Comparison of reinforcement learning models.
SL Model Key Features Pros Cons
1 Q-Learning Model-free, TD-based, can handle large state spaces Simple to implement, can handle complex decision-making tasks Can be slow to converge, requires careful selection of learning rate and exploration rate
2 SARSA Model-free, TD-based, can handle large state spaces Can handle dynamic and stochastic environments Can be slow to converge, requires careful selection of learning rate and exploration rate
3 Deep Q-Networks (DQN) Uses deep neural networks to approximate Q-values, can handle high-dimensional state spaces Can handle complex environments with high-dimensional state spaces Can be unstable during training, requires careful selection of hyperparameters
4 Policy Gradients Model-free, policy-based, can handle continuous action spaces Can handle complex environments with continuous action spaces Can be slow to converge, requires careful selection of learning rate and exploration rate
5 Actor-Critic Model-free, combines value-based and policy-based methods, can handle continuous action spaces Can handle complex environments with continuous action spaces Can be slow to converge, requires careful selection of learning rate and exploration rate
6 Proximal Policy Optimization (PPO) Model-free, policy-based, can handle continuous action spaces Can handle complex environments with continuous action spaces, more stable than policy gradients Can be slow to converge, requires careful selection of learning rate and exploration rate
7 Monte Carlo Tree Search (MCTS) Model-based, tree-based search, can handle large state spaces Can handle complex decision-making tasks, can make decisions based on incomplete information Computationally expensive, may not work well in continuous state or action spaces
4 Agriculture: An Overview
Agriculture is the cultivation of crops, livestock, and other natural resources for human consumption or use. It has been an integral part of human civilization since ancient times. The major branches of agriculture, including agronomy, horticulture, forestry, livestock, fisheries, agriculture engineering, and economics.
Agriculture tasks can be categorized into three main sub-areas: pre-harvesting, harvesting, and post-harvesting. Fig. 6 provides a detailed depiction of these sub-domains of agricultural tasks. During pre-harvesting tasks, the selection of suitable crops for the climate and soil conditions of the area is focused on by farmers. The land is prepared, seeds are sown, and the crops are maintained through techniques like irrigation, pesticide use, pruning, and other practices. Estimating the yield is also an important aspect of pre-harvesting activities, involving activities such as yield mapping and fruit counting to predict production and make appropriate arrangements for harvesting or post-harvesting. The process of collecting mature crops or fruits from the fields, known as harvesting, is focused on ensuring that the crops meet market quality standards. Post-harvesting activities, which include storage and processing systems, play a crucial role in maintaining the quality of the harvested produce. In each stage of farming, important factors such as soil fertility, irrigation, and pest control need to be considered.
Figure 6: Overview of the stages and steps involved in agriculture.
5 Applications of Machine Learning in Agriculture
ML techniques have gained significant popularity in the field of agriculture, offering promising solutions to enhance farming practices, optimize crop yield, and minimize losses. These techniques can be effectively utilized at various stages of agriculture, as demonstrated in the following sections.
5.1 Machine Learning Applications in Pre-Harvesting
Numerous studies have explored the application of ML in pre-harvesting. Radhika et al. [41] compared support vector regression (SVR) and multi-layer perceptron (MLP) for atmospheric temperature prediction, highlighting the superior performance of SVR. Srunitha et al. [42] demonstrated the effectiveness of support vector machines (SVM) in accurately classifying non-sandy soils, suggesting the potential for ML to enhance soil property prediction. Nahvi et al. [43] focused on improving the accuracy of estimating soil temperature using the Extreme Learning Machine (ELM) technique with the Self-adaptive evolutionary (SaE) algorithm. Akshay et al. [44] emphasized the importance of precise soil moisture measurement and management for optimal crop yield, proposing the K-Nearest Neighbors (K-NN) algorithm for data analysis. Uzal et al. [45] found that Convolutional Neural Networks (CNNs) outperformed traditional methods for estimating the number of seeds per soybean pod, demonstrating robust behavior against seasonal changes. Reddy et al. [46] proposed an IoT-based smart irrigation system that predicts water requirements using a decision tree algorithm to enhance irrigation efficiency. AI et al. [47] focused on using CNNs for automated crop disease identification, achieving high accuracy and developing a WeChat applet for disease recognition. Mridha et al. [48] proposed a deep learning-based method for early detection and identification of plant diseases, outperforming traditional models and offering a web-based AI tool. Muhammed et al. [49] presented a user-friendly crop recommendation system (UACR) that leverages AIoT architecture to assist farmers in selecting suitable crops based on location, rain level, and soil type.
5.2 Machine Learning Applications in Harvesting
Significant advancements have been made in utilizing ML for harvesting purposes. Durmucs-Bunlu et al. [50] discussed the development of a mobile autonomous robot designed for agricultural operations, with the goal of enhancing production efficiency through precision farming, disease diagnosis, and soil analysis. De Baerdemaeker et al. [51] introduced Octinion’s fully autonomous picking robot for strawberry cultivation, addressing labor shortages in the industry. This robot effectively identifies and harvests ripe strawberries, ensuring their quality while neatly arranging them in punnets. Zhang et al. [52] presented an improved deep learning-based method for tomato ripeness classification in a harvesting robot. The study explored various techniques for dataset augmentation to enhance accuracy using convolutional neural networks (CNN). Kestur et al. [53] proposed MangoNet, a deep learning system that utilizes a convolutional neural network and contour-based object detection for mango detection and counting in RGB images, enabling estimation of mango yield. Altaheri et al. [54] described the development, testing, and validation of SWEEPER, an autonomous robot for harvesting sweet pepper fruit in greenhouses. SWEEPER employs an industrial arm and advanced sensing technology to perform its harvesting tasks. Furthermore, Bogue et al. [55] provided insights into recent commercial developments in fruit picking robots, focusing specifically on strawberry and other soft fruit harvesting, while examining their commercial prospects.
5.3 Machine Learning Applications in Post-Harvesting
In the field of post-harvesting, ML has demonstrated its potential in various applications. Al-Hilali et al. [56] address the challenges faced by Saudi Arabian growers by proposing a visionary solution that harnesses the power of computer vision and RGB images. Their groundbreaking system extracts quality features and achieves precise categorization of dates with an impressive 80% accuracy using a back propagation neural network. Purandare et al. [57] present an end-to-end system aimed at reducing post-harvest losses in agriculture. This system provides real-time data, suggestions for optimal harvesting, and predictive analytics for warehouse management. By employing statistical and probabilistic techniques, it mitigates storage and transportation losses while supporting informed decision-making. Chakravarthy et al. [58] introduce an automated system for classifying mangoes based on their condition using image processing techniques. Leveraging MATLAB and the k-means clustering algorithm, this study extracts features from mango images and evaluates the accuracy of the classification process, ultimately enhancing efficiency in fruit grading and classification. Another notable application is the tomato grading machine vision system described by Ireri et al. [59]. This system utilizes RGB images to achieve accurate and reliable grading by detecting calyx and stalk scars, identifying defected regions, and recognizing various grading categories based on color and texture features. With its high accuracy and potential for inline sorting, this system plays a crucial role in maintaining quality standards during tomato production. Duysak et al. [60] employ the K-nearest neighbor (KNN) algorithm and stepped frequency continuous wave radar (SFCWR) to accurately determine grain quantity in a silo. Through the analysis of back-scattering signals, range profiles, and feature combinations, they achieve an impressive $96.71\%$ accuracy, highlighting the effectiveness of ML in assessing grain quantity with precision. Lutz et al. [61] propose a hierarchical model called Boosted Continuous Non-spatial whole Attribute Extraction (BCNAE) for detecting post-harvest losses in sekai-ichi apples. This model comprises three levels: manual separation, shape fitting, and feature refinement through image segmentation. The BCNAE framework demonstrates effective detection and reduction of post-harvest losses, offering precise solutions in apple production.
5.4 Machine Learning Application in Disease Detection
Researchers have extensively utilized ML and image processing techniques to identify diseases in various types of plants during pre-harvesting, harvesting and post-harvesting. By employing these methods, they aim to accurately detect and diagnose plant diseases, enabling timely intervention and effective management strategies.
An automated approach for leaf disease detection in agriculture is proposed in [62] for tomato plants. The model aims to detect diseases at early stages, enhancing crop production. SVM, Random Forest, and Logistic Regression algorithms are compared, with SVM outperforming the others. The study lacks specific details on dataset size and evaluation metrics. Similarly, a ML and computer vision-based apple disease detection system is proposed in [63]. It achieves $98.63\%$ accuracy by segmenting diseased regions and utilizing features like DWT and color histograms.
For accurate and efficient multi-class plant disease detection, a deep learning object detection model is proposed in [64], surpassing the performance of existing models with a mean average precision of $91.2\%$ and an F1-score of $95.9\%$. Another study [65] introduces a deep convolutional neural network (DCNN) for early diagnosis of apple tree leaf diseases (ATLDs). The proposed DCNN achieves $98.82\%$ accuracy, showing fast convergence, fewer parameters, and high robustness. In [66], an accurate identification approach for common apple leaf diseases is proposed using a deep convolutional neural network. It achieves a $97.62\%$ accuracy with reduced parameters. A lightweight convolutional neural network, RegNet, is introduced in [67] for rapid and accurate identification of apple leaf diseases, surpassing other pre-trained models with a $99.23\%$ overall accuracy. The research in [68] presents a CNN model with $19$ layers for accurate classification of Marsonina Coronaria and Apple Scab diseases in apple leaves, achieving $99.2\%$ accuracy on a dataset of $50,000$ images. Another ML-based approach for automated detection of apple tree diseases is proposed in [69], achieving high accuracy of approximately $95\%$ for disease classification. A deep convolutional neural network model is proposed in [70] for real-time detection of apple leaf diseases, achieving an overall accuracy of $99.31\%$ and surpassing other models. The research in [71] utilizes fine-tuned pre-trained CNN models for accurate apple crop disease detection, with DenseNet201 achieving $98.75\%$ accuracy.
In [72], an automated software solution using image processing techniques is proposed for apple leaf disease detection, based on color and texture features. Finally, an algorithm for accurate segmentation of diseased parts in apple leaf images is proposed in [73], achieving an accuracy of $96.4\%$ with the $k$ nearest neighbor classifier. Table LABEL:tab:long provides a comprehensive overview of the diverse applications of ML in the field of agriculture.
Table 5: Application of ML at various stages of agriculture.
Sn Application Year Features Models Pros Cons Ref.
Pre-harvesting
1 Soil Classification 1995 The author is trying to give a quantitative soil classification. Nueral Network The effectiveness of this technique is due to its ability to classify soil quantitatively. Requires a large amount of high-quality training data to achieve accurate results [74]
2 Weather prediction 2009 This paper describes the use of SVM for predicting weather. SVM The finding that SVM consistently outperformed MLP suggests that SVM may be a more effective approach for weather prediction tasks. The selection of appropriate parameters is critical for the successful implementation of SVM in weather predictio. [41]
3 Soil Classification 2016 This paper explains SVM based classification of the soil types. Soil SVM Model is showing less accuracy SVM is a powerful ML technique that can handle large and complex datasets. SVM models are sensitive to the choice of kernel function and its associated parameters, which requires careful tuning to optimize performance. [42]
4 Soil Temperature 2016 Soil in this paper the measurement is for soil temperature (ST) by AI is presented. ELM, SaE-ELM, genetic p rogramming (GP) and ANN Higher generalization performance and prediction accuracy compared to traditional algorithms. SVM-based soil classification shows lower accuracy compared to alternative classification methods. [43]
5 Soyabeen pods 2018 38 types of variety of soyabeen pods has been classified CNN, SVM CNNs offer a more accurate and efficient approach for feature extraction compared to traditional methods. Training and fine-tuning CNNs can be computationally expensive, especially for large datasets. [45]
6 Seed test 2019 This paper proposes a method for classifying plant seedlings using a dataset consisting of around 5,000 images representing 960 unique plants from 12 species at different developmental stages. CNN The proposed solution can assist farmers in optimizing their crop production, which can potentially lead to higher yields and profits. Both CNN and SVM models require careful selection of hyperparameters, which is time-consuming process and require expert knowledge [75]
7 Irrigation plaining 2019 The proposed solution will be developed by establishing a distributed wireless sensor network, wherein each region of the farm would be covered by various sensor modules which will be transmitting data on a common server. IoT, SVR, RF, Wireless sendors Sutaiable and computational efficient approach by Iot and ML Reliability concerns due to the possibility of sensor failures or connectivity issues. [76]
8 Crop predicor and weed detector 2020 The aim of the paper is to provide farmers with a user-friendly and straightforward application that can assist them in selecting the appropriate crops for their land and detecting and eliminating weeds in their crop fields. CNN, WSN, naive bayes Precise management of the farm sector can lead to increased productivity and cost savings for farmers. The model require significant initial investment and ongoing maintenance costs for hardware, software, and data collection [77]
9 Seed test 2020 A mobile application has been developed that quickly detects and classifies seed images with high accuracy using CNN, one of the deep learning techniques. CNN, Inceptionv3, Xception and Inception ResNetV2 (i) The mobile application can detect and classify seed images with high accuracy using CNN, a powerful deep learning technique. (ii) The use of advanced models such as Inceptionv3, Xception, and InceptionResNetV2 can enhance the performance of the application. The system not able to identify all seed species accurately, especially rare or less common ones.. [78]
10 Irrigation plaining 2020 The authur has propsed a irrigation system using ML CNN , DT (i) The use of CNN architecture helps in capturing complex patterns and relationships in the data. (ii) The decision tree algorithm helps in improving the interpretability and explainability of the model The model’s performance may be affected by the quality and accuracy of the data used for training, as well as changes in weather conditions [79]
11 Irrigation planing 2020 This paper presents an advanced technology based smart system to predict the irrigation requirements of a field by sensing of ground parameter like moisture of the soil, temperature- humidity and water level using ML algorithm The paper proposes a smart system that uses ML algorithms to predict the irrigation requirements of a field by sensing ground parameters such as soil moisture, temperature-humidity, and water level. RF (i) The system offers a solution to the over fitting problem faced by existing algorithms like K-mean and SVM. (ii) The system has the capacity to realize a fully automated irrigation scheme. The system may not perform optimally under extreme weather conditions or in areas with highly variable soil types [44]
12 Irrigation planing 2020 This research presents a smart irrigation system that utilizes ML algorithms to predict the water requirements of a crop. DT Better yield prediction. Decision trees are prone to over-fitting the training data [46]
13 Disease detection 2020 CNN is used to automatically identify crop diseases Inception-ResNet-v2 This model for automatic crop disease detection save time and reduce human effort. Shows lower accuracy when compared with other ML model [47]
14 Soil 2020 The study aims to develop a model that can predict soil fertility indices for different villages based on soil reaction, organic carbon, and nutrient content. ELM with different activation functions like sine-squared, Gaussian radial basis etc The approach is cost-effective and efficient, as it eliminates the need for expensive and time-consuming laboratory tests. Not able to capture the complex interactions between different soil parameters and their impact on soil fertility indices [80]
15 Crop recommendation system 2021 The paper presents a yield prediction system for farmers that is both practical and user-friendly. SVM, ANN, RF, multi linear regression These models handles large amount of data and complexity of data The effectiveness of crop selection assistance may depend on the availability and accuracy of data on soil and climate conditions [81]
16 disease detection and croprecommendation system 2021 This paper presents location and season-based crop recommendation system CNN i.e. Sequential and VGG-16. The recommendation feature can also help farmers to diversify their crops and reduce the risk of crop failure. Overfitiing when performs on limited dataset [82]
17 Seed test 2021 This paper shows that wheat seed classification. Neural network The model has High accuracy. Neural networks can be complex to design and train, requiring specialized knowledge and resources. [83]
18 Disease detection 2021 The proposed model is a hybrid system based on Convolutional Autoencoder (CAE) network and CNN for automatic plant disease detection. CNN The hybrid model can potentially improve the accuracy of disease detection compared to traditional methods. The effectiveness of the model depend on the quality and quantity of the training dataset. [84]
19 Disease detection 2021 Computer vision and AI can aid in the early detection of plant diseases, providing a solution to the challenges of continuous human monitoring and reducing the negative impacts of diseases on crops. CNN Saves time of farmer Model is dependent on high quality image data [85]
20 Crop recomendation system 2022 The author has presented an analysis of AI-driven precision farming and agriculture, including related studies, and has proposed a new crop recommendation platform that utilizes cloud-based ML. The platform is designed to aid farmers in making decisions regarding which crops to harvest by considering various known factors. KNN, DT, RF, XGBoost and SVM The platform can potentially reduce waste by providing farmers with more accurate information about when to harvest their crops, resulting in fewer spoiled or unsellable crops The platform require a significant amount of data to be effective, which may be difficult to obtain in some regions or for certain types of crops. [86]
21 Crop recomendation system 2022 The proposed technique combines IoT and AI to develop a smart agricultural system. IoT, layer Perceptron neural Multinetwork The system can optimize the use of resources such as water and fertilizers, reducing waste and environmental impact. The system require regular maintenance and updates to ensure optimal performance, which is time-consuming and expensive. [87]
22 Seed test 2022 This research work aims to explore the application of deep learning techniques for seed classification. MobileNetV2 The choice of using the MobileNetV2 model was based on its efficient memory usage and straightforward architecture. Less seed class taken [88]
23 Disease detection 2022 A model using CNN and VGG is developed to accurately classify disease-affected leaves. . CNN, VGG Classify disease-affected leaves, allowing for early detection and treatmenty Time complexity [89]
24 Disease detection 2022 The study employed a hybrid ML model to detect plant diseases and A web-based AI tool has been developed using this model for plant disease identification. CNN, AlexNet and LetNet Found it to be superior to traditional models. Does not work well in large data [48]
25 Soil Classification 2022 The study aimed to assess the effectiveness of ML modelin classifying potential acid sulfate soils in Jutland, Denmark, and also to apply a model-agnostic interpretation method on the resulting CNN model for improved understanding and insights CNN The use of CNNs for the classification of potential acid sulfate soils is a promising approach for contextual mapping in wetland areas The CNN model’s accuracy and interpretability may be influenced by the quality and quantity of training data used. [90],
26 Crop recomendation system 2023 The proposed crop recommendation system, UACR, allows farmers to input their farm location and automatically receive recommendations for the best crops to grow based on various factors such as rain level and soil type. Iot The UACR system simplifies the crop recommendation process for farmers, as they only need to input their location The absence of edge/fog nodes in some architectures may pose challenges for processing the large amounts of dat [49]
27 Irrigation plaining 2023 The research proposes an IoT-based intelligent control system for smart agriculture that utilizes sensors to collect environmental data and includes an automated irrigation system. KNN, SVM, DT, RF The system can provide real-time monitoring of environmental conditions Large computing costs. [91]
28 Intrusion detection system 2023 The paper introduces a novel system for detecting intrusions in IoT networks utilized in agriculture, using deep learning models and image-based features CNN, PSO for optimization, VGG16, Exception, and Inception model Evaluating the proposed method with different models can help identify potential weaknesses and areas for improvement. Time-consuming [92]
29 Soil Classification 2023 This study demonstrates the application of ML for soil classification using ML XGBoost,KNN imputer Model gives more accurate result Model is limited to data quality [93]
30 Pepper seeds 15 features classified based on width , length, and projected area, seed weight and densityusing ML MLP, BLR, single feature models Model has high accuracy in classification Less accuracy on more features [94]
Harvesting
31 Mobile autonomous robot for agricultural operations 2015 This technology enables accurate analysis and classification of fruit images, including those with overlapping fruits, using semantic segmentation for individual fruits and instance segmentation for overlapping fruits. R-CNN Models saves time of labour High cost of robots [50]
32 Fruit Classification 2018 We propose an efficient framework for fruit classification using deep learning. VGG-16 fine-tuned model, CNN The system can potentially reduce food waste by accurately identifying and classifying fruits, The initial cost of implementing the system high [95]
33 A fully autonomous picking robot 2018 The paper presented AI robot for fruits AI Perception Alternative for costly human picker with no damage Ensure robustness and to identify its shortcomings. [51]
34 Classifying tomato ripeness in the design of a tomato harvesting robot 2018 Improved deep learning-based classification method that improves the accuracy and scalability of tomato ripeness with a small amount of training data. CNN,DNN Less parameter calculation and higher accuracy. over-fitting probelm [52]
35 Automated fruit harvesting robot 2019 This research introduces an approach for the identification and harvesting of fruits using a robotic arm. CNN, R-CNN, VGG, YOLO The algorithm for fruit position detection and harvesting can be automated High cost of model [96]
36 RTFD real time fruit detection 2019 This study employs novel convolutional deep learning techniques based on single-shot detectors for fruit (apple and pear) detection and counting within the tree canopy. CNN, YOLO Redues fruit losses The limitation of the model in differentiating between two objects of the same category in the same grid cell may result in inaccurate fruit counting in some cases [97]
37 Method for mango detection and counting 2019 This study addresses the challenge of stable fruit detection in varying illumination conditions present in open farm environments. CNN MangoNet, a deep CNN architecture using sematic segmentation Stable fruit detection in varying illumination conditions can improve efficiency and accuracy of fruit counting and yield estimation in open farm environments This work only focused on one fruit, so the model will not perform well for other fruits. [53]
38 Fruit feature detection 2019 Use of (DCNN) to optimize features for classification and segmentation tasks in agricultural navigation VGG16 (OxfordNet),Dcnn The use of DCNNs and VGG16 for feature optimization and classification improve accuracy in agricultural navigation tasks. Models works for specific crops only [98]
39 Harvesting sweet pepper fruit in greenhouses 2019 The framework presented in this study incorporates three classification models that enable real-time classification of date fruit images based on their type, maturity, and decision regarding harvesting. AlexNet and VGG-16,CNN The model achieves best accuracy in fuits classiying The fine-tuning process of the VGG-16 model require significant computational resources and expertise to achieve optimal results. [54]
40 Fruit Yield 2019 This paper presents the deep multifaced system to increase banana fruits yield LSTM, DL The model achieved better performance Acquiring and curating such data can be time-consuming and resource-intensive. [99]
41 Fruit Detection 2020 FD should be robust to varying outdoor conditions with high signal to noise ratio. SD should help in estimating the angle at which the fruit is oriented around the stem for effective severing. Semantic Segmentation using FCN, Adaptive Image thresholding using video images using reinforced learning with decaying e-greedy algorithm FD system’s robustness and ability to handle varying outdoor conditions contribute to accurate fruit detection Computational complexity [100]
42 Fruit picking robots, 2020 The system possesses the ability to train with multiple fruit types, enabling accurate classification based on factors such as size and ripeness. Improved efficiency Multiple fruit crops support Efficiency can be improved [55]
43 Fruit picking 2020 This article presents a dual-arm aubergine harvesting robot consisting of two robotic arms configured in an anthropomorphic manner to optimize the dual workspace. Cubic SVM, water shed transformation, Point cloud extraction The use of a dual-arm configuration in the aubergine harvesting robot allows for optimized workspace Computational demands may limit real-time performance, requiring specialized hardware. [101]
44 Fruit Detectetion 2020 The system demonstrates effectiveness in detecting fruits even in environments where leaves or branches have similar colors, accommodating variability in fruit shape and irregularities in the growing environment. CNN, SVM Improved cucumber harvesting efficiency through effective fruit detection and high recognition accuracy. limited adaptability for diverse agriculture applications [102]
45 Fruit Detectetion 2020 To locate & harvest cotton balls in field conditions. Deep learning with Tiny Yolov3 with 7 conv. layers & CLoDSA based image augmentation Over $77\%$ Action Success Ratio ASR and over $97\%$ Detection Performance Future design and development research should also include alternative energy sources to decrease energy costs. [103]
46 Fruit picking 2020 The paper introduces a harvesting robot designed specifically for plucking tea leaves, utilizing Jaco 2 robotic arm Jaco 2 based leaf plucking using ProMP Increased efficiency and reduced labor in tea production. Found limitations in handling branches with varying stiffness. [104]
47 Autonomous path detection 2020 The paper presents a developing a system for autonomous path detection in semi-structured environments to enhance navigation and operation efficiency. CNN Potential for cost savings by minimizing human labor and errors. Poor lighting, occlusions, and complex surroundings can limit the accuracy of autonomous path detection. [105]
48 Fruit Detection 2020 Improved detection of small harvesting cups using gray scale-depth and color images, with practical lighting-independent applications. Faster RCNN with MobileNetV2 Accurate detection of small harvesting cups in a wide range of lighting conditions, enabling more efficient and effective operations in diverse agricultural environments. Dependency on color images for tapping position detection. [106]
49 Fruit Detection 2023 In this research, we propose lightweight fruit detection algorithm for edge CPUs, enabling 19 FPS detection speed in RTFD-CPU smartphone app.. DNN,PicoDet-S model The proposed RTFD has great potential for intelligent picking machines, Sensitivity to environmental factors that may affect fruit detection accuracy. [107]
Post-harvesting
50 Post harvesting disease 2011 The system automatically extracts quality features from external data using image processing. CNN saves time and labor cost It has both hardware and software, space taking [56]
51 Post harvesting disease 2017 The system incorporates a prediction system for warehouse managers, recommending optimal dispatch sequences and transportation conditions for minimal storage and transportation loss. IoT,ML Optimized stock dispatch, reduced loss, and real-time monitoring empower farmers for efficient crop transportation and management. Security concerns [57]
52 Post harvesting disease 2019 Image processing technique used to classify mangoes based on their condition, considering three different phases. ML; DOF; Machine Vision; Articulated Arm; Fuzzy Logic Edge Detection; RGB Eliminating the need for manual labor while preserving the fruits features and appearance. The model only takes strong features to calculate and predicts the output of the data. [58]
53 Post harvesting disease 2019 The proposed system achieved detecting calyx and stalk scars in both defected and healthy tomatoes. RBF SVM Highest accuracy The model is only for tommatoes [59]
54 Better post harvest handling 2020 Improving harvest quality and productivity for farmers through innovative solutions. Computer vision DL, YOLOV3 Assisting in the identification of healthy apples by distinguishing them from those with defects. Limitation in triaing diffrent type of data [108]
55 Better post harvest handling 2020 A ML approach utilizes radar backscattering data to accurately estimate grain quantity in silos. d KNN Easy to apply in diffrerent grains Challenges in accuaring correct radar data [60]
56 Food secuiritry 2021 A hierarchical model is proposed, integrating ontology-enabled IoT, to distinguish and separate healthy Sekai-ichi apples using Boosted Continuous Non-spatial Whole Attribute Extraction (BCNAE). IoT,ML Achieving higher accuracy and throughput while minimizing processing time. Computational complexity and accuracy-speed trade-off are potential drawbacks when aiming for higher accuracy, throughput, and speed. [109]
57 Better post harvest handling 2022 The review explores the use of sensors, IoT principles, and ML technologies in post-harvest grain monitoring, identifying potential advancements and resource utilization. IoT Enhanced grain quality monitoring Collecting sensor data requires robust measures to protect sensitive information. [61]
58 Food security 2022 The paper has investigated household-level food security RF and CHAID decision tree offers an efficient and reliable data analysis method. Prone to over-fitting The effectiveness of the model vary across different locations. [110]
59 Food security 2022 The paper has proposed detection and prediction models for identifying post-harvest deterioration zones in stored apple fruits. CNN-based U-Net, Deep-Lab , and Mask R-CNN The models enable retailers to maintain better quality control of their apple fruits, ensuring customer satisfaction. The models specifically designed for stored apple fruits only [111]
60 Food secuiritry 2022 The paper has presented Drone technology for food supply chain management . DL,ML Reduces costs associated with traditional packing and delivery methods. Drons have limited payload capcity,battery life issues [112]
61 Post harvesting disease 2023 The paper introduces a knowledge-based expert system utilizing Bayesian network for diagnosing post-harvest diseases in apple. Bayesian network Enhance the trust in the system and improve user understanding. Implementing the application for real-world field e [113]
Table 6 presents an overview of recent applications of machine learning techniques in the realm of agricultural practices, with a specific emphasis on genetic changes and plant breeding. The table chronicles research spanning from 2018 to 2023 and encapsulates various machine learning models. These models have been leveraged to accelerate crop improvement, aid in precision decision-making for speed breeding, enhance climate-resilient crop breeding, and even for automated plant phenotyping. Several references, such as [114] and [115], further underscore the significance of machine learning in advancing modern agricultural practices and outcomes.
Table 6: Overview of Recent Machine Learning Applications in Agricultural Practices, including Genetic Changes and Plant Breeding.
Application Year Model Features Ref.
Plant Breeding 2019 Nueral Network, Explainable AI The author has presented state of genomics and field phenomics, and proposes a workable path to improvement. [114]
Plant Breeding 2023 Ensemble learning, Big data, Random forest ML has used in analysing genetic data in plant breeding, accerelating the crop improvemment, developing new varities and food data sequirity. [115]
Integrating Speed Breeding 2022 Deep learning, Neural Network The author has utilized ML in speeding plant breeding and enabling precision decesion making. [116]
Climate-Resilient Smart-Crop Breeding 2022 Machine learning The author has reviewed machine learning application in crop breeding. [117]
Plant Breeding 2020 Nueral Network, SVM, RF Discuused about improving new layers of plant breeding [118]
Plant Breeding 2021 multilayer perceptron, SVM, and RF The ML models have been used for improving crop yield of soyabean [119]
Plant Breeding 2019 RF Predict in-season seed yield (SY) in soybean based on high-dimensional phenotypic trait data [120]
Plant Breeding 2019 SVM, DL The paper compares the performance of MLP and SVM methods to the Bayesian threshold genomic best linear unbiased prediction (TGBLUP) model for genomic selection in plant breeding. [121]
Estimating Phenotypic Traits 2018 Regression, DL Using color images and elevation maps, machine learning is employed to estimate emergence and biomass traits in wheat field plots. [122]
Phenotypying 2019 DL, CNN CNN based modeling is used in computer vision for automated plant phenotyping, enabling high-throughput analysis of plant characteristics. [123]
6 Statistical and Bibliometric Analysis
In this section, we present the results of the statistical analysis conducted to comprehend the extent of research activity in the field of ML for agriculture. Statistical data has been gathered, revealing a substantial growth trend in recent years, highlighting it as one of the most dynamic and vibrant research domains.
Figure 7: The amount of research carried out in machine learning for agriculture. Data source: Scopus on May 19, 2023.
Fig. 7 showcases the volume of research conducted in the intersection of machine learning (ML) and agriculture. The search criteria encompassed the terms ‘machine learning’ and ‘agriculture’ in the title, abstract, or keywords of the publications. Each bar signifies a distinct year, with its height denoting the count of research articles during that year. A perusal of the bars highlights the evolving trajectory of research in this domain. A marked surge is evident between 2018 and 2020. An exponential trendline is plotted, boasting an $R^{2}$ value of $0.989$. This coefficient of determination indicates how well the trendline fits the actual data. An $R^{2}$ value nearing 1, as seen here, suggests that the trendline provides an exceptional fit, underscoring a robust correlation between years and publication count. The near-perfect fit of the exponential trendline, accompanied by the high $R^{2}$ value, reiterates the burgeoning research momentum in ML applied to agriculture. The bar chart succinctly captures this escalating research enthusiasm over the years, highlighting the field’s rising prominence.
Figure 8: Research contribution by countries in machine learning for agriculture. Data source: Scopus on May 19, 2023.
Fig. 8 presents the research contribution of different countries in ML for agriculture. The pie diagram visually represents the distribution of research publications, with each slice corresponding to a specific country or the ‘Others’ category. The size of each slice represents the proportionate contribution of that country to the total number of publications. India has the largest slice, indicating the highest number of research publications among the listed countries, followed closely by the United States and China. This pie diagram demonstrates a clear overview of the relative contributions of different countries in the field of ML for agriculture.
Table 7: Publication distribution by document Type. Data source: Scopus on May 16, 2023.
Document Type No. of Publications Percentage (%)
Article 3190 49.63
Conference Paper 2378 37.00
Review 362 5.63
Book Chapter 286 4.45
Conference Review 147 2.29
Book 28 0.44
Data Paper 23 0.36
Editorial 13 0.20
Total 6427 100
In addition, Table 7 provides the distribution of publications by document type in the field of ML for agriculture. Moreover, the table provides the percentage distribution of each document type, which offers a relative perspective on their contribution to the total number of publications. For instance, articles accounted for $49.63\%$ of the total publications, while conference papers constituted $37\%$. The remaining document types had smaller percentages, ranging from $0.20\%$ for editorials to $4.45\%$ for book chapters. In summary, the table offers a concise overview of the publication distribution in ML for agriculture, shedding light on the prevalence of various document types and their respective proportions within the field.
Figure 9: Number of publications by different types of publishers. Data source: Scopus on May 19, 2023.
(a) Publication in IEEE
(b) Publication in Elsevier
(c) Publication in Springer
(d) Publication in MDPI
(e) Publication in miscellineous
(f) Publication in conferences
Figure 10: Distribution of articles on the intersection of ML and agriculture across various journals.
The contribution of publishers is depicted in Fig. 9. The majority of papers are published by Elsevier, followed by Mdpi, Springer Nature, Wiley, IEEE, and others. More detailed statistics for publication house are provided in Fig. 10. This statistical analysis provides valuable insights into the volume of research conducted and the distribution of publications in ML for agriculture. The substantial growth trend observed in recent years reinforces the dynamic and vibrant nature of this research domain.
Figure 11: The bibliometric analysis of keywords in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of keywords (Fig. 11) in ML application in agriculture provides valuable insights into the key areas of focus within this field. In this paper, the open-source application VoSviewer [124] has been used for bibliometric analysis. By examining the frequency and trends of specific keywords, researchers can identify the most prominent topics and research directions. The analysis reveals that keywords such as ‘precision farming’, ‘crop yield prediction’, ‘plant disease detection’, and ‘soil analysis’ are among the most frequently mentioned in research publications. This indicates a strong emphasis on utilizing ML techniques for optimizing farming practices, improving crop productivity, and addressing agricultural challenges. By understanding the prevalent keywords, researchers and practitioners can gain a better understanding of the areas where ML is making significant contributions and guide future research efforts.
Figure 12: The bibliometric analysis of countries in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of countries (Fig. 12) in ML application in agriculture offers a comprehensive view of the research contributions from different nations. By assessing the number of research publications, researchers can identify the countries that are leading in this field. The analysis reveals that India, the United States, and China are among the top contributors in terms of research publications. These countries demonstrate a strong commitment to leveraging ML techniques for agricultural applications and have made substantial contributions to the advancement of the field. Additionally, the analysis highlights the contributions of other countries collectively categorized as ”Others,” indicating a global interest in ML application in agriculture. By recognizing the leading countries, researchers can foster collaborations and knowledge-sharing initiatives to further accelerate the progress in ML for agricultural innovation.
Figure 13: The bibliometric analysis of organisations in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of organisations in ML application in agriculture is shown in Fig. 13. The advancement of scientific research relies on the collaborative efforts of numerous institutions worldwide. Aarhus University, Addis Ababa Science and Technology University, and Agricultural Research Center are just a few examples of the organizations contributing to the knowledge base. These institutions have produced a significant number of research documents and garnered citations, demonstrating their impact on their respective fields. With a strong link strength, they have fostered connections and collaborations within the scientific community, further enhancing the progress of scientific inquiry.
7 Case Studies and Examples
The use of AI in agriculture has enormous potential to revolutionize the way farming is conducted, from improving crop management to reducing costs and simplifying tasks for farmers. This potential is being realized by a number of start-ups and applications that leverage AI to address these challenges and improve productivity and efficiency in the agricultural sector. Here are some examples of how AI is being practically utilized in agriculture:
  • Kisan GPT: It, launched on March 15, 2023, is an AI chatbot that caters to the underserved agriculture sector of India. It provides guidance on various topics such as irrigation, pest control, and crop cultivation. Unlike other popular AI chatbots where users are required to type their queries, Kisan GPT offers the option of voice input using a microphone. It supports 10 languages including Hindi, English, Gujarati, and Marathi, as well as foreign languages like Japanese, Portuguese, and Indonesian. With India being an agricultural powerhouse, this AI chatbot could be a game-changer for the agriculture sector. Farmer-friendly and accessible, it bridges the information gap between farmers and experts, aiding decision-making and enhancing crop management. With AI-powered chatbots, like Kisan GPT, farmers can achieve greater efficiency and make informed choices, benefiting India’s agricultural prowess [125].
  • Jiva: Jiva provides a range of services, including micro-financing and AI-based systems, empowering farmers to unlock the potential of AI. With Jiva, farmers can enhance crop health, streamline operations, and increase profitability, ushering in a promising future for agriculture. AI-powered systems like Jiva offer personalized recommendations for optimal farming practices, resulting in higher crop yields and improved profitability [126].
  • See and Spray: The Blue River Technology ‘See and Spray’ system uses computer vision to identify individual plants and selectively apply herbicides, reducing the amount of chemicals needed and improving crop yields [127].
  • FarmX app: FarmX is an AI-powered app that helps farmers optimize their irrigation and fertilizer usage. The app uses ML algorithms to analyze data from various sources, such as soil moisture sensors, weather data, and satellite imagery. With this information, it generates recommendations for farmers on when and how much water and fertilizer to apply to their crops. The app can also provide alerts if there are any issues with irrigation, such as leaks or blockages. By optimizing irrigation and fertilizer usage, the FarmX app helps farmers improve crop yields while reducing water and fertilizer waste.
AI and robotics are revolutionizing the agricultural industry and helping farmers to address the challenges they face. As technology continues to advance, we can expect even more innovative solutions to emerge in the future.
8 Top 10 AI and Robotics Companies/Start-Up Disrupting Agriculture
The use of AI and robotics in agriculture has been gaining momentum in recent years, and some companies have emerged as global leaders in this field. The top 10 global companies using AI and robotics in agriculture, based on their valuation, are Benson Hill Biosystems, Iron Ox, FarmWise, Carbon Robotics, CropIn, Blue River Technology, Aerobotics, Prospera, AgroScout, and Root AI. The details about this company is provided in Table 8. These companies have developed various technologies, including ML, computer vision, and robotics, to optimize crop growth, monitor plant health, and detect diseases early. By leveraging these technologies, they are helping farmers make data-driven decisions and achieve higher yields while reducing their environmental footprint. Overall, the emergence of these companies is a testament to how AI and robotics can revolutionize agriculture and contribute to a more sustainable and efficient food system.
Table 8: Top 10 global Companies Using AI and Robotics in Agriculture based on their valuation.
Rank Name of Company Headquarters Founded Year Service Website
1 Benson Hill Biosystems USA 2012 Uses ML for analyzing crop gene influence https://bensonhillbio.com/
2 Iron Ox USA 2015 Uses robotics and AI to ensure optimal plant growth https://www.ironox.com
3 FarmWise USA 2016 Builds adaptive robots to farm individual plants https://farmwise.io/
4 Carbon Robotics USA 2018 Makes advanced robotics accessible to improve crop yield https://carbonrobotics.com/
5 CropIn India 2010 Helps with remote sensing, weather advisory, scheduling and monitoring farm activities, educating farmers, and monitoring crop health http://www.cropin.com/
6 Blue River Technology USA 2011 Uses computer vision and robotics to identify lettuce plants with precision http://bluerivertechnology.com
7 Aerobotics South Africa 2014 Uses aerial imagery and ML to help farmers identify pests and diseases early https://www.aerobotics.com/
8 Prospera Israel 2014 Develops computer vision technologies to monitor and analyze plant health and provide actionable insights to growers https://www.prospera.ag/
9 AgroScout Israel 2017 Uses data from multiple sensors, ML, and deep learning to detect and monitor diseases in the field http://agro-scout.com/
10 Root AI USA 2017 Aims to create intelligent robots to help growers build the farms of the future https://root-ai.com/
9 Challenges and Limitations of Machine Learning in Agriculture
While ML and IoT applications [128] in smart farming are changing the way work is done in the agricultural sector, there are still some limitations to the implementation of AI in agriculture that require further research. Few challenges and limitations include as follows:
  • Subsidies and welfare schemes: This is a major issue that affects the agricultural sector in many countries. Despite the government’s efforts to provide subsidies and welfare schemes, farmers often do not receive the benefits due to bureaucratic hurdles, corruption, and lack of awareness. This has a significant impact on the livelihoods of farmers and their ability to sustain their farms [129].
  • Lack of information about water management: Water is a critical resource for agriculture, and farmers need to manage it effectively to ensure optimal crop yield. However, many farmers lack access to information about water management techniques and technologies, such as drip irrigation, water harvesting, and soil moisture sensors. This can lead to inefficient water use and reduced crop yields [130, 131].
  • Absence of agricultural marketing: Agricultural marketing is essential for farmers to sell their produce at a fair price and reach a wider market. However, many farmers in rural areas lack access to proper markets and transportation facilities, which makes it difficult for them to sell their produce. This can result in low prices for farmers and food shortages for consumers [132, 133].
  • Lack of proper storage facilities: Proper storage facilities are essential for preserving the quality of agricultural produce and preventing post-harvest losses. However, many farmers lack access to such facilities, which can lead to spoilage and wastage of crops. This can also result in lower prices for farmers, as they are forced to sell their produce immediately after harvest [133].
  • Limited knowledge of modern agricultural practices: Agriculture is a rapidly evolving sector, with new technologies and practices being developed all the time. However, many farmers lack access to information about these new practices, which can result in low crop yields and reduced profitability. This also makes it difficult for farmers to compete in a globalized market, where other farmers may have access to the latest [134, 135].
  • Lack of awareness about ICT benefits: Information and communication technologies (ICT) can be powerful tools for farmers, allowing them to access information about weather patterns, market prices, and new technologies. However, many farmers are not aware of the benefits of these technologies or lack access to them. This can result in missed opportunities for farmers to improve their yields and profitability [135, 136].
The agricultural industry and farmers are deeply concerned with sustainability, and any implementation of AI technology must be capable of not only meeting but exceeding existing sustainability standards [137]. However, the use of AI may also lead to less sustainable practices and outcomes if not implemented properly. One potential consequence of relying on AI in agriculture is the possibility of an increase in pesticide usage, potentially leading to the adoption of more potent and hazardous pesticides without human oversight. Furthermore, if human labor is substituted with large, weighty robots, this could exacerbate preexisting issues linked to soil compaction, a problem already associated with the use of heavy machinery in farming [138].
The safe deployment of AI poses a significant challenge that cannot be overlooked. Farmers have raised concerns that the use of AI-powered robots [139] and drones in agriculturecould result in increased risks and harm to the environment, as there may be less attention given to potential harm to humans. This underscores the importance of carefully considering all potential consequences of AI implementation in farming and taking appropriate measures to mitigate any negative impacts.
Trust building within the agricultural sector is quite challenging as well [140]. When autonomous machines are deployed on farms and collect data that is transmitted back to agribusinesses, it becomes challenging to establish trust, particularly since farmers typically have limited knowledge about how their data is being utilized. In the agricultural industry, data privacy represents a significant concern, as farmers may need to compromise their privacy to reap the benefits and potential advantages of AI technology on their farms. Consequently, there is a growing apprehension about the possibility of AI infringing on farmers’ privacy [141].
Additionally, the ownership and control of the data retrieved, stored, and used by the AI system is another issue that needs to be addressed. Some other challenges and limitations of AI in agriculture include: U
  • Limited data availability: Limited data availability for training AI models, due to the lack of technology infrastructure and limited connectivity in rural areas, is a major challenge for implementing AI in agriculture [142].
  • Crop and soil variability: Crop and soil variability, which makes it challenging to create a ‘unique’ model that works for all conditions, requires AI models to be tailored to specific local conditions.
  • Weather variability: Weather variability, which makes it difficult to predict crop yields and adapt to changing conditions, requires AI models to be constantly updated and refined to account for changing weather patterns.
  • High cost of technology: The high cost of technology required for implementing AI in agriculture, such as sensors, drones, and other hardware, can be a barrier to adoption for small-scale farmers.
  • Lack of technical expertise: The lack of technical expertise among farmers and rural communities can make it difficult to implement and maintain AI technologies in agriculture.
  • Limited awareness: Limited awareness among farmers about new technologies and ethical concerns surrounding AI in agriculture is another barrier to adoption.
While AI has the potential to revolutionize the agricultural sector, there are still many challenges and limitations that need to be addressed before widespread adoption can occur. Further research is needed to overcome these challenges and ensure that the benefits of AI in agriculture are fully realized.
10 Future Directions of Machine Learning in Agriculture
AI and AI-driven solutions are touted as the panacea of our age, and agriculture is one area where the benefits of AI applications are pragmatic and realistic. Here are some future directions of ML in agriculture:
  • IoT standardization: It is crucial to develop credibility and establish a market for a new concept, but the integration of physical objects into the internet poses several challenges regarding the compatibility of existing internet protocols and applications with these objects [143].
  • IoT data: The reliability, consistency, and volume of agricultural data may be compromised due to equipment malfunctions, network node failures, post-processing errors, or pest and disease outbreaks. These data gaps can negatively impact the accuracy of calculations and hinder the effectiveness of IoT applications in agriculture [144].
  • Regulatory issues: It is imperative to establish clear regulatory and legal frameworks for the management and ownership of farm data, particularly with regards to the relationship between farm laborers and data companies. These regulations may vary from country to country, taking into account factors such as service delivery, technological complexity, market competition, and data security. This variance in regulatory practices across different regions can impact the application of IoT in specific agricultural contexts, such as monitoring and agro-food supply chains [136].
  • Security: Smart agriculture relies on IoT devices to operate effectively, but these devices are at risk of physical tampering that could impede their function. Common sources of tampering include theft, damage by livestock and rodents, relocation, and connectivity problems. Such tampering can result in a range of issues, including data loss, equipment damage, and increased expenses for farmers. As such, it is important to take steps to secure IoT devices and minimize their vulnerability to physical interference [145, 146].
  • Market issues: Given the agricultural industry’s tendency to operate with slim profit margins, it is important to carefully strategize the integration of IoT-enabled technologies by considering their potential long-term benefits [144, 147].
  • Drones and unmanned aerial vehicles: Unmanned aerial vehicles (UAVs)[148] offer a cost-effective and broad approach to environmental monitoring by capturing images and collecting data about specific scenes. As practical applications for drone technology continue to advance, the use of drones in agriculture is expected to expand in the coming years [149].
  • Driverless tractors: Robotic agriculture is on the horizon and is expected to materialize in the next decade or so. In this future, autonomous tractors will replace human labor and perform all aspects of farming. These tractors will be fitted with sensors that will allow them to monitor obstacles and apply necessary inputs. Additionally, they will be able to perform required practices, ultimately leading to increased efficiency and productivity in the agriculture sector [147].
  • Automated irrigation systems: Automated irrigation systems, when combined with historical weather data, can accurately predict the necessary resources required for optimal crop growth. These systems utilize real-time ML to maintain the desired soil conditions, which leads to an increase in average yields. One of the most significant benefits of these systems is the reduction in labor required for irrigation. Additionally, they have the potential to reduce production costs. The system’s reliance on historical weather data ensures that resources are used efficiently, which helps to reduce waste. Automated irrigation systems provide farmers with greater control over the irrigation process, resulting in healthier crops and higher yields [147].
  • Crop health monitoring: Conventional crop health monitoring techniques are laborious and largely categorical. However, companies are now competing to develop hyperspectral imaging and 3D laser scanning technologies that can enhance precision and accuracy by collecting vast amounts of data. By using deep learning ML techniques, an alert-based system can be developed to improve crop protection [147].
11 Conclusion
The agriculture sector continues to be a crucial priority, supporting the livelihoods of around 50% of the global population. As the potential for the adoption of AI in agriculture becomes increasingly apparent, the gap between current agricultural practices and AI adoption is rapidly closing. In this study, we have conducted an exhaustive review of ML applications in agriculture, providing a comprehensive understanding of the intersection between ML and agriculture.
To establish the foundations, we have explored the basic principles of ML and its relevance to agriculture. In addition, we have presented case studies that exemplify practical applications of AI in the agricultural domain, showcasing how it is used to address real-world challenges and drive significant changes. Furthermore, we have examined the top ten agro-tech companies, highlighting their innovative use of AI and ML to solve pressing issues faced by farmers and drive transformative advancements in the industry. Despite the promising potential, the adoption of AI in agriculture comes with certain challenges and limitations. We have discussed these challenges and limitations in detail, shedding light on the obstacles that need to be overcome for a successful transition. Looking ahead, we have projected the future directions of ML in agriculture through a series of point-wise projections. By considering global demand, infrastructure availability and affordability, as well as national and international initiatives, we have outlined the fundamental changes that are already underway and driving the transformation of the agricultural sector.
Inevitably, the future of agriculture belongs to early adopters of AI. As the gap between current practices and the full potential of AI in agriculture continues to shrink, it becomes evident that embracing AI technologies is essential for sustained progress and success in the industry. By being at the forefront of AI adoption, these early adopters will play a crucial role in shaping the future of agriculture and ensuring its continued growth, sustainability, and prosperity.
Acknowledgments
The authors would like to express their gratitude to the Ministry of Human Resource Development (MHRD) of India for providing financial assistance to the first author through grant number MHR01-23-200-428, and the Council of Scientific & Industrial Research (CSIR) of India for supporting the second author with grant number 09/143(0989)/2019-EMR-I during this research.